# HodgeFormer: Transformers for Learnable Operators on Triangular Meshes through Data-Driven Hodge Matrices

### v3

@@ 2 Related Work
Recent advances in geometric deep learning have adapted convolutional neural networks from 2D to 3D geometry for classification and segmentation tasks. These methods adapt convolution operations for sampled 3D manifolds by exploiting relationships between mesh elements, i.e. vertices, edges and faces, to construct convolutional filters.
MeshCNN, introduced by Hanocka et al. [11], proposed edge-based mesh convolution, mesh pooling and mesh unpooling operations. However, the method was limited by the fixed topology requirement and potential loss of geometric information during pooling. DualConvMesh-Net [31] attempted to bridge local and global features through geodesic and Euclidean convolutions, but faces computational challenges with large meshes.
@@
Another category of methods utilizes subdivision schemes to enable convolution operations on meshes. SubdivNet [7] and Subdivision-based Mesh Convolutional Networks [12] enforce regular structure through remeshing or subdivision techniques. The latter defines convolution operations directly on 3D triangle meshes via subdivision structure, treating mesh faces as fundamental convolution units analogous to 2D pixels. However, these subdivision-driven approaches suffer from several limitations, potentially introducing geometric artifacts, failing to preserve fine local features, requiring specific connectivity patterns, such as loop subdivision sequences, and incurring high computational costs from remeshing operations.
@@
Spectral methods are based on the eigen-decomposition of discrete operators, typically the Laplace-Beltrami operator, utilizing the translation, rotation, and scale invariance of operations in the spectral domain. HodgeNet [33] introduces a learnable class of sparse operators on meshes, built from standard constructions in discrete exterior calculus. It employs standard geometry processing operations and Hodge star operators ($\star_{0}$ and $\star_{1}$) to encode per-vertex and per-edge features. Spectral approaches by [8] leverage the graph Laplacian for convolution, with subsequent works like CayleyNets [17] improving spectral filtering design. However, these methods struggle with computational complexity due to the high computational cost of eigen-decomposition. To avoid this limitation, Surface Networks [15] leverage Graph Neural Networks (GNNs) to learn polynomial expansions of both intrinsic (Laplacian) and extrinsic (Dirac) operators to capture mesh geometry, enabling them to process irregular mesh structures directly.
Shape descriptors built on the Laplace-Beltrami operator have been particularly successful in capturing geometric properties. The Heat Kernel Signature (HKS) [35] captures multiscale geometric information through the heat diffusion process, providing an isometry-invariant point descriptor. The Wave Kernel Signature (WKS) [1] offers an alternative based on quantum mechanical principles, achieving better feature localization. Building on these descriptors, several neural architectures have emerged. DiffusionNet [32] leverages diffusion-based features to achieve robust segmentation and classification on partial and non-isometric shapes. Geodesic CNNs [25] introduce geometry-adapted convolutional and pooling layers that operate in the spectral domain, utilizing HKS and WKS features as input channels for learning intrinsic shape representations.
Other methods like MeshWalker [16] use random walks for mesh exploration. Instead of trying to adapt CNNs to irregular mesh structures, they represent meshes through random walks on the mesh surface. Each walk explores both local geometry and global topology of the mesh, employing an RNN-based architecture to process the features along the walks. Despite being able to work well with small datasets, MeshWalker requires long walks for large meshes, which can lead to runtime and memory issues.
LaplacianNet [29] introduces a deep learning framework for 3D meshes encoding mesh connectivity using Laplacian spectral analysis. The network uses Mesh Pooling Blocks (MPBs) to split the mesh surface into local pooling patches and aggregate both local and global features. A mesh hierarchy is built from fine to coarse using Laplacian spectral clustering, making the method robust to different triangulations and isometric transformations. The method is dependent on eigen-decomposition and spectral clustering.
@@
Recent Transformer adaptations display potential in capturing global relationships in 3D meshes. The Laplacian Mesh Transformer [18] applies Transformer layers to Laplacian eigenvectors and geometric features. MeshMAE [21] introduces self-supervised learning through masked autoencoders for mesh understanding. MeshFormer [24] addresses scalability by combining mesh simplification with hierarchical graph-based attention mechanisms. Mesh Transformer (MeT) [36] proposes a two-stream architecture that processes triangle and cluster features in parallel, using Laplacian eigenvectors for positional encoding. Despite their effectiveness, these Transformer-based methods face significant computational challenges. The eigen-decomposition required for Laplacian features and the quadratic complexity of attention mechanisms limit their applicability to large-scale meshes, while the need for complex architectural components increases implementation and training difficulty.
@@ 3.2 Transformers
A Transformer is a function $T:\mathbb{R}^{N\times d_{m}}\to\mathbb{R}^{N\times d_{m}}$, where $d_{m}$ is the embedding dimension defined by the composition of $L$ Transformer layers $T_{1}(\cdot),\dots,T_{L}(\cdot)$ as follows:
@@
The function $f_{l}(\cdot)$ transforms each feature independently of the others and is usually implemented with a small two-layer feedforward network. $A_{l}(\cdot)$ is the self-attention function and is the only part of the Transformer that acts across sequences. The self-attention function $A_{l}(\cdot)$ computes, for each position, a weighted sum of all positions in the sequence. Formally, for input $x\in\mathbb{R}^{N\times d_{m}}$.
@@
where $W_{Q},W_{K}\in\mathbb{R}^{d_{m}\times d_{k}}$ , $W_{V}\in\mathbb{R}^{d_{m}\times d_{v}}$. Typically, $d_{k}=d_{v}=d_{m}/h$, with $h$ the number of attention heads. The attention output is:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (4)
where $Q\in\mathbb{R}^{N\times d_{k}}$, $K\in\mathbb{R}^{N\times d_{k}}$, $V\in\mathbb{R}^{N\times d_{v}}$, and the output shape is $\mathbb{R}^{N\times d_{v}}$. For multi-head attention, this process is repeated in parallel for each head, and the outputs are concatenated and projected back to $d_{m}$. The feedforward network is applied independently to each position (row) and is typically a two-layer MLP.
@@
Figure 3: Multi-head Hodge Attention applied on latent vertex features $x_{v}$. The multi-head attention mechanism learns data-driven Hodge Star matrices $\star_{0}^{-1}$ and $\star_{1}$.
@@ 4.1 HodgeFormer Layer
We propose a novel deep learning architecture shown in Fig. 2, named “HodgeFormer”, that leverages the multi-head attention mechanism for learning data-driven Hodge Laplacian operators on k-forms based on the parametrization of the Hodge Star $\star_{k}$. If $x_{k}\in\mathbb{R}^{n_{k}\times d_{k}}$ is a k-form on mesh elements with $k\in\{v,e,f\}$, respectively, of size $n_{k}$ and $W_{Q_{k}}\in\mathbb{R}^{d_{k}\times d_{k}}$, $W_{K_{k}}\in\mathbb{R}^{d_{k}\times d_{k}}$, and $W_{V_{k}}\in\mathbb{R}^{d_{k}\times d_{k}}$ are the learnable projection matrices transforming $x_{k}$ to corresponding representations $Q_{k}$, $K_{k}$, and $V_{k}$ as in Eq. 4, then we can parametrize the Hodge Star operator $\star_{k}$ as:
$\star_{k}(x_{k})$ $=\sigma\left({\frac{Q_{k}K_{k}^{T}}{\sqrt{d_{k}}}}\right),$ $\text{for }k\in\{v,e,f\}$ (5)
where $\sigma\left(\cdot\right)$ is the row-wise softmax non-linearity. For the inverse Hodge Star operators, i.e., $\star_{k}^{-1}$, we employ dedicated linear maps ${Q_{k}},{K_{k}}$. We define the HodgeFormer Layer $H_{l}$ as the function that operates and updates k-forms on vertices, edges and faces:
@@
$x_{v},x_{e},x_{f}\enspace:=$ $\enspace A_{H_{l}}(x_{v},x_{e},x_{f})$ (7)
$x_{v}=A_{v}(x_{v},x_{e})\enspace=$ $\enspace L_{v}(x_{v},x_{e})\cdot V_{v}(x_{v})$ (8)
$x_{e}=A_{e}(x_{v},x_{e},x_{f})\enspace=$ $\enspace L_{e}(x_{v},x_{e},x_{f})\cdot V_{e}(x_{e})$ (9)
$x_{f}=A_{f}(x_{e},x_{f})\enspace=$ $\enspace L_{f}(x_{e},x_{f})\cdot V_{f}(x_{f})$ (10)
and $G_{l}(\cdot)$ a function that transforms the values of each k-form independently and is implemented with separate two-layer feedforward networks with residual connections. The Hodge Laplacian $L_{v}$, $L_{e}$, $L_{f}$ for vertices, (0-form), edges (1-form) and faces (2-form) can be defined as :
$\begin{split}L_{v}\enspace:=\enspace&\,\star_{0}^{-1}(x_{v})\cdot d_{0}^{T}\cdot\star_{1}(x_{e})\cdot d_{0}\\ L_{e}\enspace:=\enspace&\,d_{0}\cdot\star_{0}^{-1}(x_{v})\cdot d_{0}^{T}\cdot\star_{1}(x_{e})\,+\\ &\quad\,\,\,\,\star_{1}^{-1}(x_{e})\cdot d_{1}^{T}\cdot\star_{2}(x_{f})\cdot d_{1}\\ L_{f}\enspace:=\enspace&\,d_{1}\cdot\star_{1}^{-1}(x_{e})\cdot d_{1}^{T}\cdot\star_{2}(x_{f})\end{split}$ (11)
@@ 4.2 End-to-end Architecture
The input consists of features extracted from the mesh, along with the sparse oriented incidence matrices $d_{0}$ and $d_{1}$. Dedicated embedding layers operate separately on vertex, edge, and face features and map them to target latent dimensions $d_{v}$, $d_{e}$, and $d_{f}$. Then, the latent embeddings for each mesh element are transformed by a mixture of sequentially placed HodgeFormer and Transformer layers and updated accordingly. The final embeddings are passed to the task head, which maps them to the task-specific output dimensions. Putting together HodgeFormer layers with vanilla Transformer layers can be thought of as combining local operators with global mixers. Different mixing strategies of HodgeFormer and vanilla Transformer layers can be applied.
@@ 5.1 Evaluation Setup
The HodgeFormer architecture is implemented in PyTorch [28] with standard backpropgation. The BFS operations for local neighborhood extraction are based on sparse matrix operations using the graphblas framework [5], [6]. All experiments and measurements are performed on a single Nvidia RTX 4090 GPU with 24GB VRAM, paired with Intel i9-14900KF processors and 64GB of system RAM.
@@ 5.2 Experiments and Results
Method Type Acts Eigen SHREC11 Cube Human COSEG COSEG COSEG
On Decomp. (split-10) Engrav. Simplified Vases Chairs Aliens
MeshCNN [11] cnn e No 91.0% 92.2% 85.4% 92.4% 93.0% 96.3%
PD-MeshNet [26] cnn ef No 99.1% 94.4% 85.6% 95.4% 97.2% 98.2%
MeshWalker [16] rnn v No 97.1% 98.6% n/a 99.6% 98.7% 99.1%
HodgeNet [33] mlp v Yes 94.7% n/a 85.0% 90.3% 95.7% 96.0%
DiffusionNet [32] mlp v Yes 99.5% n/a 90.8% n/a n/a n/a
LaplacianNet [29] mlp v Yes n/a n/a n/a 92.2% 94.2% 93,9%
Laplacian2Mesh [10] cnn v Yes 100.0% 91.5% 88.3% 94.6% 96.6% 95,0%
SubDivNet [12] cnn f No 99.5% 98.9% 91.7% 96.7% 96.7% 97.3%
MeT [36] trns f Yes n/a n/a n/a 99.8% 98.9% 99.3%
HodgeFormer (ours) trns vef No 98.7% 95.3% 90.3% 94.3% 98.8% 98.3%
Table 3: Performance of different methods on mesh classification and mesh segmentation tasks. Mesh classification is evaluated on the 30-class SHREC11[20] dataset evaluated on splits of 10 samples per class (split-10) and the Cube Engraving dataset [11]. Mesh segmentation is evaluated on the Human dataset in its simplified version of Milano et al. [26] with hard ground truth labels at faces, as well on the Shape COSEG dataset for the categories of Vases, Chairs and Aliens. HodgeFormer achieves results comparable to the state-of-the-art without spectral features, eigenvalue decomposition operations or complex complementary structures. We report their base architecture, the mesh elements they operate on (v: vertices, e: edges, f: faces) and whether they depend on eigen-decomposition methods.
For the mesh segmentation task, our goal is to predict labels for every face of a mesh. We evaluate HodgeFormer architecture on four datasets: on the Human-part-segmentation dataset in its simplified version [26] with hard ground truth labels at faces, and the Vases, Chairs, and Aliens categories from the Shape COSEG dataset. In all experiments, we optimize against a class-weighted cross-entropy loss with an additional label smoothing factor of $0.2$. For each dataset, we report the highest performing variant of our architecture. For Human and Vases datasets, our highest performing model was a $6$-layered HodgeFormer model mixed with 2 Transformer layers in a $2:1$ ratio, whereas for Chairs and Aliens datasets, our highest performing model was a $6$-layered HodgeFormer model followed by 2 Transformer layers. As Tab. 3 demonstrates, our approach performs comparatively to state of the art without spectral features, eigenvalue decomposition operations, or complex complementary structures.

### v4

@@ 2 Related Work
Mesh convolutions. Recent advances in geometric deep learning have adapted convolution operations for sampled 3D manifolds by exploiting relationships between mesh elements, i.e. vertices, edges and faces, to construct convolutional filters. MeshCNN, introduced by Hanocka et al. [11], proposed edge-based mesh convolution, mesh pooling and mesh unpooling operations. However, the method was limited by the fixed topology requirement and potential loss of geometric information during pooling. DualConvMesh-Net [34] attempted to bridge local and global features through geodesic and Euclidean convolutions, but faces computational challenges with large meshes. Another category of methods utilizes subdivision schemes to enable convolution operations on meshes. SubdivNet [7] and Subdivision-based Mesh Convolutional Networks [12] enforce regular structure through remeshing or subdivision techniques. The latter defines convolution operations directly on 3D triangle meshes via subdivision structure, treating mesh faces as fundamental convolution units analogous to 2D pixels. However, these subdivision-driven approaches suffer from several limitations, potentially introducing geometric artifacts, failing to preserve fine local features, requiring specific connectivity patterns, such as loop subdivision sequences, and incurring high computational costs from remeshing operations.
@@
Spectral methods are based on the eigen-decomposition of discrete operators, typically the Laplace-Beltrami operator, utilizing the translation, rotation, and scale invariance of operations in the spectral domain. HodgeNet [36] introduces a learnable class of sparse operators on meshes, built from standard constructions in discrete exterior calculus. It employs standard geometry processing operations and Hodge star operators ($\star_{0}$ and $\star_{1}$) to encode per-vertex and per-edge features. Spectral approaches by [8] leverage the graph Laplacian for convolution, with subsequent works like CayleyNets [17] improving spectral filtering design. However, these methods struggle with computational complexity due to the high computational cost of eigen-decomposition. To avoid this limitation, Surface Networks [15] leverage Graph Neural Networks (GNNs) to learn polynomial expansions of both intrinsic (Laplacian) and extrinsic (Dirac) operators to capture mesh geometry, enabling them to process irregular mesh structures directly. Shape descriptors built on the Laplace-Beltrami operator have been particularly successful in capturing geometric properties. The Heat Kernel Signature (HKS) [38] captures multiscale geometric information through the heat diffusion process, providing an isometry-invariant point descriptor. The Wave Kernel Signature (WKS) [1] offers an alternative based on quantum mechanical principles, achieving better feature localization. Building on these descriptors, several neural architectures have emerged. DiffusionNet [35] leverages diffusion-based features to achieve robust segmentation and classification on partial and non-isometric shapes. Geodesic CNNs [26] introduce geometry-adapted convolutional and pooling layers that operate in the spectral domain, utilizing HKS and WKS features as input channels for learning intrinsic shape representations. LaplacianNet [31] introduces a deep learning framework for 3D meshes encoding mesh connectivity using Laplacian spectral analysis. The network uses Mesh Pooling Blocks (MPBs) to split the mesh surface into local pooling patches and aggregate both local and global features. A mesh hierarchy is built from fine to coarse using Laplacian spectral clustering, making the method robust to different triangulations and isometric transformations. The method is dependent on eigen-decomposition and spectral clustering.
Graph Learning-based E(n)-Equivariant Graph Neural Networks (EGNN) [33] efficiently extend message passing to geometric graphs by maintaining scalar invariant features and equivariant coordinates for each node, using relative distances to preserve rotations, translations, and reflections. Equivariant Mesh Neural Networks (EMNN)[39] build on this by adding surface-aware messages that leverage mesh face information. EMNN further incorporates multiple vector channels and hierarchical pooling to capture long-range dependencies while maintaining E(3)-equivariance.
@@
Transformer adaptations display potential in capturing global relationships in 3D meshes. The Laplacian Mesh Transformer [18] applies Transformer layers to Laplacian eigenvectors and geometric features. MeshMAE [21] introduces self-supervised learning through masked autoencoders for mesh understanding. MeshFormer [24] addresses scalability by combining mesh simplification with hierarchical graph-based attention mechanisms. Mesh Transformer (MeT) [40] proposes a two-stream architecture that processes triangle and cluster features in parallel, using Laplacian eigenvectors for positional encoding. Despite their effectiveness, these Transformer-based methods face significant computational challenges. The eigen-decomposition required for Laplacian features and the quadratic complexity of attention mechanisms limit their applicability to large-scale meshes, while the need for complex architectural components increases implementation and training difficulty.
@@ 3.2 Transformers
A Transformer is a function $T:\mathbb{R}^{n\times d_{m}}\to\mathbb{R}^{n\times d_{m}}$ where $d_{m}$ is the embedding (model) dimension, defined by the composition of $L$ Transformer layers $T_{1}(\cdot),\dots,T_{L}(\cdot)$:
@@
Here, $f_{l}(\cdot)$ transforms each row of shape $\mathbb{R}^{1\times d_{m}}$ independently of the others and is usually implemented with a small two-layer feedforward network. $A_{l}(\cdot)$ is the self-attention function and is the only component acting across sequence positions. The self-attention function $A_{l}(\cdot)$ computes, for each position, a weighted sum of all positions in the sequence. Formally, for input $x\in\mathbb{R}^{n\times d_{m}}$:
@@
Typically $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d_{m}\times d_{h}}$ where $d_{h}=d_{m}/{h}$, with $h$ the number of attention heads, and $Q\in\mathbb{R}^{n\times d_{h}}$, $K\in\mathbb{R}^{n\times d_{h}}$, $V\in\mathbb{R}^{n\times d_{h}}$. The attention output is:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{h}}}\right)V$ (4)
and the output shape is $\mathbb{R}^{n\times d_{h}}$. For multi-head attention, this process is repeated in parallel for each head; the outputs are concatenated and projected back to $d_{m}$. The feedforward network is applied independently to each position (row) and is typically a two-layer MLP.
@@
Figure 3: Multi-head Hodge Attention applied on latent vertex features $x_{v}$. The multi-head attention mechanism learns data-driven Hodge Star matrices $\star_{0}^{-1}$ and $\star_{1}$. Layer Norm (LN) is applied on $Q$, $K$ matrices to increase training stability.
@@ 4.1 HodgeFormer Layer
We propose a novel deep learning architecture shown in Fig. 2, named “HodgeFormer”, that leverages the multi-head attention mechanism for learning data-driven Hodge Laplacian operators on k-forms based on the parametrization of the Hodge Star $\star_{k}$. Let $x_{k}\in\mathbb{R}^{n_{k}\times d_{k}}$ a k-form on mesh elements with $k\in\{v,e,f\}$, respectively, of size $n_{k}$ and let $W_{Q_{k}}\in\mathbb{R}^{d_{k}\times d_{h_{k}}}$, $W_{K_{k}}\in\mathbb{R}^{d_{k}\times d_{h_{k}}}$, and $W_{V_{k}}\in\mathbb{R}^{d_{k}\times d_{h_{k}}}$ the learnable projection matrices transforming $x_{k}$ to corresponding representations $Q_{k}$, $K_{k}$, and $V_{k}$ as in Eq. 4. For simplicity we set $d_{k}=d$ for $k\in\{v,e,f\}$, and $d_{h_{k}}=d_{h}=d/h$ for $h$ attention heads. Then we can parametrize the Hodge Star operator $\star_{k}$ as:
$\star_{k}(x_{k})$ $=\sigma\left({\frac{Q_{k}K_{k}^{T}}{\sqrt{d_{h}}}}\right),$ $\text{for }k\in\{v,e,f\}$ (5)
where $\sigma\left(\cdot\right)$ is the row-wise softmax non-linearity. For the inverse Hodge Star operators, i.e., $\star_{k}^{-1}$, we employ dedicated linear maps ${Q_{k}},{K_{k}}$.
We define the HodgeFormer Layer $H_{l}$ as the function that operates and updates k-forms on vertices, edges and faces:
@@
$x_{v},x_{e},x_{f}\kern 5.0pt:=$ $\kern 5.0ptA_{H_{l}}(x_{v},x_{e},x_{f})$ (7)
$x_{v}=A_{v}(x_{v},x_{e})\kern 5.0pt=$ $\kern 5.0ptL_{v}(x_{v},x_{e})\cdot V_{v}(x_{v})$ (8)
$x_{e}=A_{e}(x_{v},x_{e},x_{f})\kern 5.0pt=$ $\kern 5.0ptL_{e}(x_{v},x_{e},x_{f})\cdot V_{e}(x_{e})$ (9)
$x_{f}=A_{f}(x_{e},x_{f})\kern 5.0pt=$ $\kern 5.0ptL_{f}(x_{e},x_{f})\cdot V_{f}(x_{f})$ (10)
and $G_{l}(\cdot)$ a function that transforms the values of each k-form independently and is implemented with separate two-layer feedforward networks with residual connections.
From Eq. 1, the Hodge Laplacians $L_{v}$, $L_{e}$, $L_{f}$ are defined for vertices (0-form), edges (1-form) and faces (2-form) as :
$\begin{split}L_{v}\kern 5.0pt:=\kern 5.0pt&\,\star_{0}^{-1}(x_{v})\cdot d_{0}^{T}\cdot\star_{1}(x_{e})\cdot d_{0}\\ L_{e}\kern 5.0pt:=\kern 5.0pt&\,d_{0}\cdot\star_{0}^{-1}(x_{v})\cdot d_{0}^{T}\cdot\star_{1}(x_{e})\,+\\ &\quad\,\,\,\,\star_{1}^{-1}(x_{e})\cdot d_{1}^{T}\cdot\star_{2}(x_{f})\cdot d_{1}\\ L_{f}\kern 5.0pt:=\kern 5.0pt&\,d_{1}\cdot\star_{1}^{-1}(x_{e})\cdot d_{1}^{T}\cdot\star_{2}(x_{f})\end{split}$ (11)
The HodgeFormer layer $H_{l}$ extends the standard transformer layer to independently operate and update $k$-form features using multi-head attention via mesh-based Hodge Laplacians and Stars, closely matching the residual and feedforward update principles in canonical transformer networks.
@@ 4.2 End-to-end Architecture
The input consists of features extracted from the mesh, along with the sparse oriented incidence matrices $d_{0}$ and $d_{1}$. Dedicated embedding layers, discussed in Sec. 4.5, operate separately on vertex, edge, and face features and map them to target latent dimensions $d_{v}$, $d_{e}$, and $d_{f}$ (Sec. 4.5). Then, the latent embeddings for each mesh element are transformed by a mixture of sequentially placed HodgeFormer and Transformer layers and updated accordingly. The final embeddings are passed to the task head, which maps them to the task-specific output dimensions. Putting together HodgeFormer layers with vanilla Transformer layers can be thought of as combining local operators with global mixers. Different mixing strategies of HodgeFormer and vanilla Transformer layers can be applied.
@@ 4.5 Embedding Layer
Method Type Acts Eigen SHREC11 Cube Human COSEG COSEG COSEG
On Decomp. (split-10) Engrav. Simplified Vases Chairs Aliens
HodgeNet [36] mlp v Yes 94.7% n/a 85.0% 90.3% 95.7% 96.0%
DiffusionNet [35] mlp v Yes 99.5% n/a 90.8% n/a n/a n/a
LaplacianNet [31] mlp v Yes n/a n/a n/a 92.2% 94.2% 93,9%
Laplacian2Mesh [10] cnn v Yes 100.0% 91.5% 88.6% 94.6% 96.6% 95,0%
MeT [40] trns f Yes n/a n/a n/a 99.8% 98.9% 99.3%
MeshCNN [11] cnn e No 91.0% 92.2% 85.4% 92.4% 93.0% 96.3%
PD-MeshNet [27] cnn ef No 99.1% 94.4% 85.6% 95.4% 97.2% 98.2%
MeshWalker [16] rnn v No 97.1% 98.6% n/a 99.6% 98.7% 99.1%
SubDivNet [12] cnn f No 99.5% 98.9% 91.7% 96.7% 96.7% 97.3%
EMNN (MC+H) [39] gnn ef No 100% n/a 88.7% n/a n/a n/a
EGNN (MC+H) [39] gnn ef No 99.6% n/a 87.2% n/a n/a n/a
HodgeFormer (ours) trns vef No 98.7% 95.3% 90.3% 94.3% 98.8% 98.3%
Table 3: Performance of different methods on mesh classification and mesh segmentation tasks. Mesh classification is evaluated on the 30-class SHREC11[20] dataset evaluated on splits of 10 samples per class (split-10) and the Cube Engraving dataset [11]. Mesh segmentation is evaluated on the Human dataset in its simplified version of Milano et al. [27] with hard ground truth labels at faces, as well on the Shape COSEG dataset for the categories of Vases, Chairs and Aliens. HodgeFormer achieves results comparable to the state-of-the-art without spectral features, eigenvalue decomposition operations or complex complementary structures. We report their base architecture, the mesh elements they operate on (v: vertices, e: edges, f: faces) and whether they depend on eigen-decomposition methods. The abbreviation ”trns” denotes the transformer architecture. Notes: The EMNN (MC+H) and EGNN (MC+H) results shown in the table are as reported by [39].
clarifies that while EGNN and EMNN [39] theoretically do not need pre-calculated Heat Kernel Signatures (HKS), the authors report precomputation of HKS values for both the SHREC and Human Segmentation benchmark datasets
@@ 5.1 Evaluation Setup
The HodgeFormer architecture is implemented in PyTorch [29] with standard backpropagation. The BFS operations for local neighborhood extraction are based on sparse matrix operations using the graphblas framework [5], [6]. All experiments and measurements are performed on a single Nvidia RTX 4090 GPU with 24GB VRAM, paired with Intel i9-14900KF processors and 64GB of system RAM.
@@ 5.2 Experiments and Results
For the mesh segmentation task, our goal is to predict labels for every face of a mesh. We evaluate HodgeFormer architecture on four datasets: on the Human-part-segmentation dataset in its simplified version [27] with hard ground truth labels at faces, and the Vases, Chairs, and Aliens categories from the Shape COSEG dataset. In all experiments, we optimize against a class-weighted cross-entropy loss with an additional label smoothing factor of $0.2$. For each dataset, we report the highest performing variant of our architecture. For the Human dataset, our highest performing model was a $6$-layered HodgeFormer model mixed with 2 Transformer layers in a $2:1$ ratio, as presented in Fig. S6a, whereas for Vases, Chairs and Aliens datasets, our highest performing model was a $6$-layered HodgeFormer model followed by two Transformer layers, as presented in Fig. S6b. As Tab. 3 demonstrates, our approach performs comparatively to state of the art without spectral features, eigenvalue decomposition operations, or complex complementary structures.
5.3 Model Variations Study
To evaluate the importance of different architecture components and parameter choices, we varied our base model in different ways and measured its performance on a segmentation task. Specifically, in Tab. 4, we compare the main architecture layers HodgeFormer and a vanilla Transformer layer along with different embedding layers, i.e., the feed-forward and the neighbor embedding layers. In Tab. 5, we test combinations of different input features based on their type by distinguishing between coordinates, normals, and areas. In Table  6, we evaluate the effect of the number of neighbors taken into account by the sparse attention component on the performance of the HodgeFormer layers. For all experiments, we consider an architecture of $4$ layers acting only on mesh vertices with embedding dimension $256$ and hidden dimension $512$, and a learning rate of $5\mathrm{e}{-4}$, and we report test accuracy results, averaged over $5$ experiment trials with random training splits. We also consider $\sqrt{n}$ neighbors for each set of $n$ mesh elements with the exception of the experiment of Tab. 6.
In Tab. 4, the HodgeFormer layers clearly outperform vanilla Transformer layers. At the same time, embedding input features while considering their neighbors offers an edge in initial training steps and results in an overall better performance. In Tab. 6, we observe the effect of the number of neighbors for vertices and edges on the HodgeFormer performance. Let $s_{k}$ for $k\in{v,e,f}$ the number of neighbors each element of a k-form attends to. For $s_{v}=1,s_{e}=1$, the performance drops significantly as the attention mechanism takes into account only self-contributions for vertices and edges. For $s_{v}=32,s_{e}=48$, which approximates $\sqrt{n}$ for meshes in the COSEG vases dataset, the performance is high. Yet, for larger numbers of neighbors, we get diminishing returns in model performance. In Tab. 5, we observe that the HodgeFormer layer is capable of achieving satisfactory results using simple coordinate features, although exhibiting lower performance and slower convergence. However, using normals as the sole feature set proves insufficient, resulting in models with poor performance and limited generalization capability.